OpenStudy (anonymous):

LaDonna has a banana. She eats 3/5 of it and then gives the rest to her younger brother, Hector. Hector eats 3/5 of what he was given and then gives the rest to his baby sister, Yolanda. Yolanda eats half of what she was given then throws the rest away. What "percentage" of the banana was thrown away? SHOW WORK

OpenStudy (anonymous):

2/25

OpenStudy (anonymous):

4%

OpenStudy (anonymous):

Sorry, 8%

OpenStudy (anonymous):

Hector got=1-3/5=2/5, Yolanda got=(2/5).(2/5)=4/25, Thrown part=(4/25).1/2=2/25=8%

OpenStudy (anonymous):

start with 100%............% wasted=100 - (3/5 + 3/5 * 3/5 + 3/5 * 3/5 * 1/2)*100